OpenStudy (anonymous):

Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

Looking for a complete answer so i understand this thanks

OpenStudy (jiteshmeghwal9):

\[x^2-9x+4\]

OpenStudy (jiteshmeghwal9):

now use the quadratic formula\[-(-9) \pm \sqrt{(-9)^2-4(1)(4)}\over2(1)\]

OpenStudy (jiteshmeghwal9):

becoz in the equation a=1 b=-9 c=4

OpenStudy (jiteshmeghwal9):

\[9 \pm \sqrt{81-16}\over2\]

OpenStudy (jiteshmeghwal9):

\[9 \pm \sqrt{65}\over2\]

OpenStudy (jiteshmeghwal9):

\[9 \pm 8.06\over2\]

OpenStudy (jiteshmeghwal9):

now make two roots one time add & one time subtract\[9+8.06\over2\]\[9-8.06\over2\]

OpenStudy (jiteshmeghwal9):

x=8.53 or 0.47 :) Understood??

OpenStudy (anonymous):

yes but i don't understand what makes this only solvable by the quadratic formula and not factoring.

OpenStudy (jiteshmeghwal9):

try to break the middle term of L.H.S such that the product becomes 4x^2 & sum becomes -9x . & u can't do this thing hence it can't be factored

OpenStudy (anonymous):

can you explain what LHS means

OpenStudy (precal):

also 4 has 1*4 2*2 as factors and you have to add them together to get 9 1+4=5 not 9 2+2=4 not 9 another reason why you have to use the quadratic formula

OpenStudy (anonymous):

thanks!

OpenStudy (precal):

anytime :)